# Weak Lensing Approximation of Wave-optics Effects from General Symmetric Lens Profiles

### v1

@@ Abstract
Gravitational wave astronomy (675); Gravitational lensing (670)
@@ 1 Introduction
WO effects imprint frequency-dependent amplitude and phase modulations on GW signals. Detecting these lensing-induced features enables direct measurement of lens parameters—such as the redshifted lens mass, $M_{\mathrm{Lz}}=(1+z_{\mathrm{L}})M_{\mathrm{L}}$, where $z_{\mathrm{L}}$ is the lens redshift, and the source position in the source plane—for simple point-mass and SIS lens models (Çalışkan et al., 2023a, c; Savastano et al., 2023). In contrast, identifying lensing signatures in the geometric-optics (GO) regime, corresponding to the high-frequency limit, relies on associating and interpreting multiple images of the same event (Haris et al., 2018; Hannuksela et al., 2019, 2020; Çalışkan et al., 2023b; Janquart et al., 2021a; Barsode et al., 2024; Wempe et al., 2024; Cheung et al., 2025) or so-called Type-II image distortions (Dai & Venumadhav, 2017; Ezquiaga & Zumalacárregui, 2020; Wang et al., 2021; Janquart et al., 2021b).
@@
Therefore, efficient approximations are necessary to accelerate lensing calculations, particularly in the context of population studies. Several methods exist for the high- and low-frequency limits (e.g., Takahashi, 2004; Tambalo et al., 2023a); however, these do not target the intermediate dimensionless frequency range where WO features are most prominent. Savastano et al. (2023) proposed a weak lensing approximation formulated in the time domain of the amplification factor. However, their method requires dedicated regularization and still involves the evaluation of numerical integrals.
In this Letter, we present a novel and efficient method to approximate WO effects induced by general symmetric lenses in the weak lensing regime. Section 2 reviews the general formalism of GW lensing. In Section 3, we develop the weak lensing approximation, beginning with the SIS lens in Section 3.1, and extending to general lenses in Section 3.2. In Section 3.3, we validate the approximation using the Navarro-Frenk-White (NFW) lens model. Section 4 demonstrates the consistency of our method with known approximations in the high- (Section 4.1) and low-frequency (Section 4.2) limits, while Section 5 discusses the advantages and implications of the method. We conclude with potential applications and directions for future work in Section 6.
@@ 2 Lensing of Gravitational Wave
where $\omega$ is the dimensionless frequency, $\boldsymbol{x}$ and $\boldsymbol{y}$ are the dimensionless coordinates in the lens plane and source plane, respectively. Readers are referred to Takahashi & Nakamura (2003) for detailed derivations. The integration is performed over the lens plane, with $\boldsymbol{x}$ rescaled by a characteristic scale $\xi_{0}$ of the lens. The source position $\boldsymbol{y}$ is rescaled by $\eta_{0}\equiv D_{S}\xi_{0}/D_{L}$, where $D_{S}$ and $D_{L}$ are the angular diameter distances to the source and lens, respectively. The dimensionless frequency $\omega$ is defined as Dai & Venumadhav (2017); Ezquiaga et al. (2021)
@@
where $\omega$, $\psi(x)$, and $\phi_{m}(y)$ are as previously defined, and $y$ is the magnitude of the rescaled source position. In this Letter, we focus on symmetric lenses.
@@ 3.1 Singular Isothermal Sphere
Figure 1: Comparison between the weak lensing approximation and the full analytical results for the SIS lens model. Top: Solid blue lines show the amplification factor magnitude, $|F(\omega,y)|$, as a function of the source position $y$, for $\omega=0.8$, 1.0, and 1.2 (left, middle, right), computed from the full analytical solution. Dash-dotted red lines represent the weak lensing approximation, and brown dashed lines indicate the GO limit, which is independent of $\omega$. As expected, $|F(\omega,y)|$ approaches the GO limit at large $\omega$ and $y$. Bottom: Dash-dotted blue lines show the difference between the analytical calculation and the weak lensing approximation (Analyt - WL), representing the approximation error. For comparison, dotted orange lines show the difference between the analytical result and the GO limit (Analyt - GO), characterizing the WO term amplitude. The weak lensing approximation shows excellent agreement with the analytical results, with errors consistently below 1% of the WO term amplitude.
@@
To validate this approximation, we compare it to the full analytical solution expressed in terms of Fresnel integrals (Appendix B), which is equivalent to Equation 13 without approximations (Figure 1). The approximation agrees very well with the full solution at large $\omega y^{2}/2$. Specifically, for $y>10$ and $\omega\sim 1$, the error—defined as the difference $\mathrm{Analyt}-\mathrm{WL}$”—remains below $1\%$ of the WO features, quantified by the deviation between the analytical WO solution and the GO limit ($\mathrm{Analyt}-\mathrm{GO}$). This demonstrates that the weak lensing approximation provides sufficient accuracy for astrophysical applications of WO lensing, as long as the weak lensing conditions are satisfied.
@@ 3.2 General Lenses
Therefore, it is reasonable to hypothesize that the interference term primarily arises from regions where the argument of the Bessel function is small. Noting that the integral of $xJ_{0}(x)$ yields $xJ_{1}(x)$, and that the asymptotic expansion of Bessel functions becomes valid for $x\gtrsim|\alpha^{2}-1/4|$, we focus on the region where $\omega xy$ is small, with $\alpha=1$. Motivated by this, and noting that all lens information is encoded in the lensing potential, we propose that the potential within the range 0 to $3/(4\omega y)$ provides the dominant contribution to the WO term.
For a rough but effective approximation, we model the lensing potential as a straight line between $\psi(0)$ and $\psi(3/(4\omega y))$, i.e.,
$\psi(x)\approx\left[\psi\left(\frac{3}{4\omega y}\right)-\psi(0)\right]\frac{4\omega y}{3}x.$ (28)
@@
$f(\omega y)=\left(\psi\left(\frac{3}{4\omega y}\right)-\psi(0)\right)\frac{4\omega y}{3}.$ (30)
@@ 3.3 NFW
Figure 2: Comparison between the weak lensing approximation and numerical results for the NFW lens model with $\kappa_{s}=0.5$. Top: Similar to Figure 1, but the solid blue lines show the amplification factor magnitude, $|F(\omega,y)|$, from numerical calculations performed with GRAVELAMPS (Wright & Hendry, 2022) for $\omega=0.8$, 1.0, and 1.2 (left, middle, right). Bottom: Same as the top panel, but all analytical results are replaced with numerical calculations from GRAVELAMPS, which serve as the benchmark for $|F(\omega,y)|$. For the NFW model, the WO features (Num - GO) at a given source position $y$ are less pronounced compared to the SIS model, reflecting its lower compactness. The weak lensing approximation remains accurate, with errors (Num - WL) limited to a few percent of the WO term amplitude.
@@
where $\kappa_{s}$ is the dimensionless surface density, defined as $\kappa_{s}=4\pi\rho_{s}(D_{\mathrm{L}}D_{\mathrm{LS}}/D_{\mathrm{S}})r_{s}$.
@@
Figure 2 shows the comparison for $\omega=0.8,1.0,1.2$ under the NFW model. The setup is similar to that in Figure 1. As expected, both the GO and WO effects are less prominent for the NFW model compared to the SIS case, reflecting the lower compactness of NFW halos. The weak lensing approximation agrees with the numerical results when $\omega y^{2}/2$ is large. Specifically, for $y>10$ and $\omega\sim 1$, the error—defined as the difference $\mathrm{Num-Approx}$”—is only a few percent of the WO term amplitude, quantified by the deviation between the numerical results and the GO limits ($\mathrm{Num-GO}$). Again, the weak lensing approximation remains sufficiently accurate for astrophysical applications of WO lensing, provided that the weak lensing condition is satisfied. Its successful application to the NFW model further supports the validity of the approximation.
@@ 4 Frequency Asymptotics
In the weak lensing approximation, this behavior is naturally recovered, as the approximation is valid only when $\omega y^{2}/2\gg 1$. For example, keeping $\omega y^{2}/2$ large and fixed, a small $\omega\ll 1$ requires $y\gg 1$ to maintain the condition. According to Equation 31, and noting that $\mu\propto 1/y^{2}$ and $\Delta_{(1)}\propto 1/y^{4}$ for the NFW model, both the bGO and WO terms vanish as $\omega y^{3}\gg 1$, even at fixed $\omega y^{2}$. In contrast, the QGO approximation yields a WO term with amplitude $4\kappa_{s}/(\omega y^{2})^{2}$, which remains finite for fixed $\omega y^{2}$ even as $\omega\to 0$.
Figure 4: Comparison between the weak lensing approximation, QGO approximation, and numerical results for the NFW lens model with $\kappa_{s}=0.5$ and $\omega=0.1$. The layout of the top and bottom panels follows Figure 2, but here the brown dashed line shows the amplification factor magnitude, $|F(\omega,y)|$, from the QGO approximation (Takahashi, 2004). An inset in the top panel highlights the range $4\leq y\leq 8$. The weak lensing approximation outperforms the QGO approximation at low $\omega$, particularly at small $y$, and provides an accurate description once $\omega y^{2}/2\gtrsim 5$.
The weak lensing approximation also performs well at low $\omega$, before reaching the fully unperturbed limit. Figure 4 shows the comparison for $\omega=0.1$ under the NFW model, with both the QGO and weak lensing approximations compared to the numerical results. The QGO approximation becomes inaccurate at low $\omega$, diverging for $y<5$. In contrast, the weak lensing approximation remains stable and agrees well with the numerical results even when only $\omega y^{2}/2\sim 5$ is reached. This demonstrates the robustness of the weak lensing approximation for lensing by low-mass structures.
@@ 5 Discussion
Gravitational lensing in the weak lensing regime is particularly important for realistic GW detections, as most standard sirens at high redshift are expected to be affected by weak lensing. However, the weak lensing signal is dominated by the collective effect of numerous dark matter halos at large impact parameters. Even when neglecting interference between lenses and simply summing individual contributions, obtaining the amplification factor from populations of general lenses remains computationally expensive using numerical methods. For each amplification factor, it requires evaluating complex families of contours in the time domain (Villarrubia-Rojo et al., 2024). Although Savastano et al. (2023) use the weak lensing condition to simplify the contours, their method still involves one-dimensional numerical integration for each approximated contour. In contrast, our analytical weak lensing approximation requires no numerical integrations and is significantly faster.
Another major advantage of our method is that most numerical approaches evaluate the amplification factor in the time domain first, including the approach where a weak lensing approximation is implemented. However, transforming back to the frequency domain is nontrivial. Efficient computation requires Fast Fourier Transform techniques with adaptive resolution, and careful regularization is necessary to avoid artifacts in the frequency-domain results (Tambalo et al., 2023a, b; Villarrubia-Rojo et al., 2024). These challenges become particularly severe at large $y$ or $\omega$. In contrast, our method operates entirely in the frequency domain, avoiding the need for adaptive resolution and notoriously annoying regularizations, and is thus free from these complications.
Although the weak lensing approximation appears simple, it offers valuable physical insights. The approximation of the WO term involves only the lensing potential at $x=0$ and $x=3/(4\omega y)$, where $x$ is the normalized impact parameter. In other words, only the inner region of the lens contributes significantly to the WO term in the weak lensing limit, and the size of this effective region decreases with increasing $y$ and $\omega$.
For symmetric lenses, the deflection angle is
@@
where $\kappa(x)=\Sigma(\xi_{0}x)/\Sigma_{\mathrm{cr}}$ is the normalized surface mass density (convergence), and $M_{<x}$ is the mass enclosed within $x$. The lensing potential is obtained by integrating the deflection angle, with the reference point chosen such that $\psi(0)=0$. In the weak lensing approximation, the potential is treated as linear within a small region, so the normalization function $f(\omega y)$ represents the average derivative of the lensing potential, proportional to the enclosed mass $M_{<3/(4\omega y)}$ divided by its critical scale $3/(4\omega y)$. The WO term then takes the form
$T_{\text{WO}}\propto\frac{M_{<\frac{3}{4\omega y}}}{y^{2}}e^{i\omega\left(\frac{1}{2}y^{2}-\phi_{\mathrm{m}}\right)}.$ (39)

### v2

@@ Abstract
Gravitational wave astronomy (675); Gravitational lensing (670); Weak gravitational lensing (1797)
@@ 1 Introduction
WO effects imprint frequency-dependent amplitude and phase modulations on GW signals. Detecting these lensing-induced features enables direct measurement of lens parameters—such as the redshifted lens mass, $M_{\mathrm{Lz}}=(1+z_{\mathrm{L}})M_{\mathrm{L}}$, where $z_{\mathrm{L}}$ is the lens redshift, and the source position in the source plane—for simple point-mass and SIS lens models (Çalışkan et al., 2023a, c; Savastano et al., 2023). In contrast, identifying lensing signatures in the geometric-optics (GO) regime, corresponding to the high-frequency limit, relies on associating and interpreting multiple images of the same event (Haris et al., 2018; Hannuksela et al., 2019, 2020; Çalışkan et al., 2023b; Janquart et al., 2021a; Barsode et al., 2025; Wempe et al., 2024; Cheung et al., 2025) or so-called Type-II image distortions (Dai & Venumadhav, 2017; Ezquiaga & Zumalacárregui, 2020; Wang et al., 2021; Janquart et al., 2021b).
@@
Therefore, efficient approximations are necessary to accelerate lensing calculations, particularly in the context of population studies. Several methods exist for the high- and low-frequency limits (e.g., Takahashi, 2004; Choi et al., 2021; Tambalo et al., 2023a; Yarimoto & Oguri, 2025); however, these do not target the intermediate dimensionless frequency range where WO features are most prominent. Savastano et al. (2023) proposed a weak lensing approximation formulated in the time domain of the amplification factor. However, their method requires dedicated regularization and still involves the evaluation of numerical integrals.
In this paper, we present a novel and efficient method to approximate WO effects induced by general symmetric lenses in the weak lensing regime. Section 2 reviews the general formalism of GW lensing. In Section 3, we develop the weak lensing approximation, beginning with the SIS lens in Section 3.1, and extending to general lenses in Section 3.2. In Section 3.3, we validate the approximation using the Navarro-Frenk-White (NFW) lens model. Section 4 demonstrates the consistency of our method with known approximations in the high- (Section 4.1) and low-frequency (Section 4.2) limits, while Section 5 discusses the advantages and implications of the method. We conclude with potential applications and directions for future work in Section 6.
@@ 2 Lensing of Gravitational Wave
where $\omega$ is the dimensionless frequency, $\boldsymbol{x}$ and $\boldsymbol{y}$ are the dimensionless coordinates in the lens plane and source plane, respectively. Readers are referred to Takahashi & Nakamura (2003) for detailed derivations. The integration is performed over the lens plane, with $\boldsymbol{x}$ rescaled by a characteristic scale $\xi_{0}$ of the lens. The source position $\boldsymbol{y}$ is rescaled by $\eta_{0}\equiv D_{S}\xi_{0}/D_{L}$, where $D_{S}$ and $D_{L}$ are the angular diameter distances to the source and lens, respectively. The dimensionless frequency $\omega$ is defined as (Dai & Venumadhav, 2017; Ezquiaga et al., 2021)
@@
where $\omega$, $\psi(x)$, and $\phi_{m}(y)$ are as previously defined, and $y$ is the magnitude of the rescaled source position. In this paper, we focus on symmetric lenses.
@@ 3.1 Singular Isothermal Sphere
Figure 1: Comparison between the weak lensing approximation and the full analytical results for the SIS lens model. Top: Solid blue lines show the amplification factor magnitude, $|F(\omega,y)|$, as a function of the source position $y$, for $\omega=0.8$, 1.0, and 1.2 (left, middle, right), computed from the full analytical solution. Dash-dotted red lines represent the weak lensing approximation, and brown dashed lines indicate the GO limit, which is independent of $\omega$. As expected, $|F(\omega,y)|$ approaches the GO limit at large $\omega$ and $y$. Bottom: Dash-dotted blue lines show the absolute difference between the analytical calculation and the weak lensing approximation ($|\text{Analyt}-\text{WL}|$), representing the approximation error. For comparison, dotted orange lines show the absolute difference between the analytical result and the GO limit ($|\text{Analyt}-\text{GO}|$), characterizing the WO term amplitude. The weak lensing approximation shows excellent agreement with the analytical results, with errors consistently below 5% of the WO term amplitude for $y>10$.
@@
To validate this approximation, we compare it to the full analytical solution expressed in terms of Fresnel integrals (Appendix B), which is equivalent to Equation 13 without approximations (Figure 1). The approximation agrees very well with the full solution at large $\omega y^{2}/2$. Specifically, for $y>10$ and $\omega\sim 1$, the error—defined as the absolute difference $|\mathrm{Analyt}-\mathrm{WL}|$”—remains below $5\%$ of the WO features, quantified by the absolute deviation between the analytical WO solution and the GO limit ($|\mathrm{Analyt}-\mathrm{GO}|$). This demonstrates that the weak lensing approximation provides sufficient accuracy for astrophysical applications of WO lensing, as long as the weak lensing conditions are satisfied.
@@ 3.2 General Lenses
Therefore, it is reasonable to hypothesize that the interference term primarily arises from regions where the argument of the Bessel function is small. Noting that the integral of $xJ_{0}(x)$ yields $xJ_{1}(x)$, and that the asymptotic expansion of Bessel functions becomes valid for $x\lesssim\sqrt{\alpha+1}$, we focus on the region where $\omega xy$ is small, with $\alpha=1$. Motivated by this, and noting that all lens information is encoded in the lensing potential, we propose that the potential within the range 0 to $\sqrt{2}/(\omega y)$ provides the dominant contribution to the WO term.
For a rough but effective approximation, we model the lensing potential as a straight line between $\psi(0)$ and $\psi(\sqrt{2}/(\omega y))$, i.e.,
$\psi(x)\approx\left[\psi\left(\frac{\sqrt{2}}{\omega y}\right)-\psi(0)\right]\frac{\omega y}{\sqrt{2}}x.$ (28)
@@
$f(\omega y)=\left(\psi\left(\frac{\sqrt{2}}{\omega y}\right)-\psi(0)\right)\frac{\omega y}{\sqrt{2}}.$ (30)
@@ 3.3 NFW
Figure 2: Comparison between the weak lensing approximation and numerical results for the NFW lens model with $\kappa_{s}=0.5$. Top: Similar to Figure 1, but the solid blue lines show the amplification factor magnitude, $|F(\omega,y)|$, from numerical calculations performed with GRAVELAMPS (Wright & Hendry, 2022) for $\omega=0.8$, 1.0, and 1.2 (left, middle, right). Bottom: Same as the top panel, but all analytical results are replaced with numerical calculations from GRAVELAMPS, which serve as the benchmark for $|F(\omega,y)|$. For the NFW model, the WO features ($|\text{Num}-\text{GO}|$) at a given source position $y$ are less pronounced compared to the SIS model, reflecting its lower compactness. The weak lensing approximation remains accurate, with errors ($|\text{Num}-\text{WL}|$) limited to a few percent of the WO term amplitude for $y>10$.
@@
where $\kappa_{s}$ is the dimensionless surface density, defined as $\kappa_{s}=4\pi G\rho_{s}(D_{\mathrm{L}}D_{\mathrm{LS}}/D_{\mathrm{S}})r_{s}/c^{2}$.
@@
Figure 2 shows the comparison for $\omega=0.8,1.0,1.2$ under the NFW model. The setup is similar to that in Figure 1. As expected, both the GO and WO effects are less prominent for the NFW model compared to the SIS case, reflecting the lower compactness of NFW halos. The weak lensing approximation agrees with the numerical results when $\omega y^{2}/2$ is large. Specifically, for $y>10$ and $\omega\sim 1$, the error—defined as the absolute difference $|\mathrm{Num-WL}|$—is only a few percent of the WO term amplitude, quantified by the absolute deviation between the numerical results and the GO limits ($|\mathrm{Num-GO}|$). Again, the weak lensing approximation remains sufficiently accurate for astrophysical applications of WO lensing, provided that the weak lensing condition is satisfied. Its successful application to the NFW model further supports the validity of the approximation.
@@ 4.2 Low $\omega$ Limit
In the weak lensing approximation, this behavior is naturally recovered, as the approximation is valid only when $\omega y^{2}/2\gg 1$. For example, keeping $\omega y^{2}/2$ large and fixed, a small $\omega\ll 1$ requires $y\gg 1$ to maintain the condition. According to Equation 31, and noting that $\mu\propto 1/y^{2}$ and $\Delta_{(1)}\propto 1/y^{4}$ for the NFW model, both the bGO [^8] In Takahashi (2004), the bGO expansion is discussed in the high-frequency limit. However, when the variation of the lensing potential is small compared to the dimensionless frequency, as is the case in the weak lensing limit, the bGO expansion remains valid as the expansion condition is still satisfied. and WO terms vanish as $\omega y^{3}\gg 1$, even at fixed $\omega y^{2}$. In contrast, the QGO approximation yields a WO term with amplitude $4\kappa_{s}/(\omega y^{2})^{2}$, which remains finite for fixed $\omega y^{2}$ even as $\omega\to 0$. Systems with $\omega y^{2}/2\lesssim 1$ are not captured by our weak lensing approximation and may instead be treated with alternative approximations such as Choi et al. (2021).
Figure 4: Comparison between the weak lensing approximation, QGO approximation, and numerical results for the NFW lens model with $\kappa_{s}=0.5$ and $\omega=0.1$. The layout of the top and bottom panels follows Figure 2, but here the brown dashed line shows the amplification factor magnitude, $|F(\omega,y)|$, from the QGO approximation (Takahashi, 2004). An inset in the top panel highlights the range $4\leq y\leq 9$. The weak lensing approximation outperforms the QGO approximation at low $\omega$, particularly at small $y$, and provides an accurate description once $\omega y^{2}/2\gtrsim 10$.
The weak lensing approximation also performs well at low $\omega$, before reaching the fully unperturbed limit. Figure 4 shows the comparison for $\omega=0.1$ under the NFW model, with both the QGO and weak lensing approximations compared to the numerical results. The QGO approximation becomes inaccurate at low $\omega$, diverging for $y<5$. In contrast, the weak lensing approximation remains stable and agrees well with the numerical results once $\omega y^{2}/2\sim 10$ is reached. This demonstrates the robustness of the weak lensing approximation for lensing by low-mass structures.
@@ 5 Discussion
Gravitational lensing in the weak lensing regime is particularly important for realistic GW detections, as most standard sirens at high redshift are expected to be affected by weak lensing. However, the weak lensing signal is dominated by the collective effect of numerous dark matter halos at large impact parameters. Even when neglecting interference between lenses and simply summing individual contributions, obtaining the amplification factor from populations of general lenses remains computationally expensive using numerical methods. For each amplification factor, it requires evaluating complex families of contours in the time domain (Villarrubia-Rojo et al., 2025).
To address this computational challenge, several analytic approximation schemes have been proposed. Choi et al. (2021) introduced a method to approximate the amplification factor in the weak diffractive lensing regime, applicable in the low-frequency limit where $\omega y^{2}\lesssim 1$. Yarimoto & Oguri (2025) further generalized this approach beyond the strictly diffractive limit, but their formulation still assumes small $\omega$ and neglects the contribution from the central imaginary image, which is one of the most distinctive and physically important features of WO effects in the weak lensing regime and is properly captured by our approximation. On the other hand, Savastano et al. (2023) use the weak lensing condition to simplify the contours, but their method still involves one-dimensional numerical integration for each approximated contour. In contrast, our analytical weak lensing approximation requires no numerical integrations and is significantly faster.
Another major advantage of our method is that most numerical approaches evaluate the amplification factor in the time domain first, including the approach where a weak lensing approximation is implemented. However, transforming back to the frequency domain is nontrivial. Efficient computation requires Fast Fourier Transform techniques with adaptive resolution, and careful regularization is necessary to avoid artifacts in the frequency-domain results (Tambalo et al., 2023a, b; Villarrubia-Rojo et al., 2025). These challenges become particularly severe at large $y$ or $\omega$. In contrast, our method operates entirely in the frequency domain, avoiding the need for adaptive resolution and notoriously annoying regularizations, and is thus free from these complications.
Although the weak lensing approximation appears simple, it offers valuable physical insights. The approximation of the WO term involves only the lensing potential at $x=0$ and $x=\sqrt{2}/(\omega y)$, where $x$ is the normalized impact parameter. In other words, only the inner region of the lens contributes significantly to the WO term in the weak lensing limit, and the size of this effective region decreases with increasing $y$ and $\omega$.
To gain further insight into the weak lensing approximation, the connection between the WO term and the physical properties of the lens can be made explicit by expressing it in terms of the enclosed mass. For symmetric lenses, the deflection angle is
@@
where $\kappa(x)=\Sigma(\xi_{0}x)/\Sigma_{\mathrm{cr}}$ is the normalized surface mass density (convergence), and $M_{<x}$ is the mass enclosed within $x$. The lensing potential is obtained by integrating the deflection angle, with the reference point chosen such that $\psi(0)=0$. In the weak lensing approximation, the potential is treated as linear within a small region, so the normalization function $f(\omega y)$ represents the average derivative of the lensing potential, proportional to the enclosed mass $M_{<\sqrt{2}/(\omega y)}$ divided by its critical scale $\sqrt{2}/(\omega y)$. The WO term then takes the form
$T_{\text{WO}}\propto\frac{M_{<\frac{\sqrt{2}}{\omega y}}}{y^{2}}e^{i\omega\left(\frac{1}{2}y^{2}-\phi_{\mathrm{m}}\right)}.$ (39)